A square field measuring 100.0 m by 100.0 m has an area of 1.00 hectare. An acre has an area of 43,600ft2.

If a country lot has an area of 11.0 acres, what is the area in hectares?
Answer #1

Given that

1 Acre = 43600 ft^2

100m*100m = 10^4 m^2 = 1.00 hectare

Now we need to find

11.0 acres = _____ hectares

Convert acres in to ft^2

11.0 acres = (11.0 acre)*(43600 ft^2/1 Acre) = 479600 ft^2

1 ft = 0.3048 m, So

Now convert ft^2 in m^2

11.0 Acre = 479600 ft^2*(0.3048 m/1 ft)^2

11.0 Acre = 44556.3 m^2

Now convert m^2 in hectare

11.0 Acre = (44556.3 m^2)*(1 hectare/10^4 m^2)

11.0 Acre = 44556.3/10^4 = 4.45563 hectare

11.0 Acre = 4.46 hectare

Please Upvote.